OpenStudy (anonymous):

1 + cos2a= (2)/(1 + tan^2a)

OpenStudy (anonymous):

I guess that you have to prove this identity. first notice that: \(\huge1+\tan^2a=\frac{cos^2a}{cos^2a}+\frac{sin^2a}{cos^2a}=\frac{1}{cos^2a}\) so: \(\huge1+cos2a=2cos^2a\) then use formula: \(cos(2a)=cos^2a-sin^2a\) so your equatio becomes: \(\huge1+cos^2a-sin^2a=2cos^2a\) as you should know: \(sin^2a+cos^2a=1\) putting this in the last expretion gives the identity

OpenStudy (anonymous):

Which double-angle or Half-angle identity would be used?

OpenStudy (anonymous):

doublé angle

OpenStudy (anonymous):

so the one above?